# Price Experimentation and Interference

### v3

@@ 4 Revenue Management Setting
We examine a traditional revenue management scenario involving monopoly pricing. A firm or platform offers products that, at least based on observable characteristics, are deemed homogeneous and are thus priced uniformly. The firm faces an unknown to them demand function given by $D(p)$ as a function of the price $p$. The firm has a constant marginal production cost, $c\geq 0$. We make the following assumption that is kept throughout the paper and will imply Assumption 3.1 for the performance metrics we study.
@@
The function $D(p)$ is differentiable and $D^{\prime}(p)<0$, for all $p\in\mathbb{R}_{+}$. At this level of abstraction, the demand function can describe a scenario with either an infinite supply or a limited supply, where a rationing mechanism determines the consumption. In a large limit model, even with supply constraints we expect the demand function to be differentiable.
To provide a concrete example with infinite supply, the demand function can be understood as the fraction among a large population of customers buying the product at a given price. That is, $D(p)=\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}x_{i}(p)$ almost surely, where $\{x_{i}(p):i=1,...,n\}$ are i.i.d. random variables equal to one if customer $i$ buys the product at a given price $p$, and zero otherwise.
@@ 4.1 Estimands
We will consider two performance metrics of typical interest: (1) demand itself, which serves as a measure of growth, and (2) the firm’s profits (or revenue, in case costs are zero). The firm’s decision variable is price, and it is seeking to gain an understanding of how changes in price impact demand (i.e., price elasticity of demand) and subsequently affect profit. To achieve this goal, the firm is interested in estimating the difference in demand or profits generated when products are priced at a new “treatment” price relative to the demand or profits generated when products are priced at the status-quo “control” price. These differences correspond to the global treatment effects of interest that we derive below. To do so, we first define the “global” profit function as:
@@
$GTE_{D}(p)$ ::=:= $D^{\prime}(p)$ (6)
$GTE_{\pi}(p)$ ::=:= $\pi^{\prime}(p)$ (7)
$=$ $D(p)+(p-c)D^{\prime}(p)$
@@ 4.2 Experiment: Design, Model, and Estimator
Given that the demand function is unknown, we assume the firm uses price experimentation to estimate the $GTE$s of interest. As mentioned above, canonical A/B-tests can randomize at either side of the market, and our model does not preclude either possibility. That said, we do note that when conducting price experiments in practice, companies prefer to refrain from randomizing prices on the customer side to adhere to the basic principle that each customer should be offered the same price for the same product (Phillips, 2021).
@@
which correspond to the demand for control and treated subjects scaled by the control and treatment fractions, respectively. Henceforth, we refer to $D_{0}$ and $D_{1}$ as the control and treatment demand functions. Again, at this level of abstraction these demand functions could, for example, incorporate supply constraints and rationing mechanisms.
Recall that in a customer-side experiment treatment and control customers will observe all products priced at the treatment and control prices, respectively, with the proportion of customers assigned to the treatment group being $q$. In contrast, in a product-side experiment, a customer will face an array of products, a fraction $q$ of which are priced at $p_{1}$ and the rest at $p_{0}$.
We make the following assumption that implies Assumption 3.2 and that is kept throughout the paper. {assumption} The functions $D_{0}(p_{0},p_{1},q)$, and $D_{1}(p_{0},p_{1},q)$ are differentiable for all $p_{0},p_{1}\in\mathbb{R}_{+}$ and $10<q<$ 1. In addition, $p_{0},p_{1}\geq c$. Note that the assumption precludes the case of perfect substitution between control and treatment products. Under perfect substitution, a negligible undercut of the treatment relative to the control price would result in a discrete increase of the treatment demand; as a result, the demand function would not be continuous. Even though, as initially stated, the firm is selling a single type of product, we can rationalize the smoothness assumption in at least one of two ways: (1) products can be identical in terms of observable characteristics but there may be latent horizontal differentiation; (2) there could be search frictions that do not allow customers to always buy the lowest-priced product.
We introduce the following additional assumption on experimental demand functions that implies Assumption 3.2. {assumption} The following holds:
@@
The first two assumptions are natural as they imply that control and treatment products are (imperfect) substitutes. The third assumption states that if the control and treatment prices are the same, then both the scaled control and treatment demand functions are equivalent to the market demand function; this assumption is implied by constant returns to scale. This assumption is satisfied by demand systems for which the control and treatment demand quantities are proportional to their shares in the marketplace. Indeed, this assumption will be satisfied for the model we later study in Section 7, as well for a number of other common demand systems.
@@
$\pi_{0}(p_{0},p_{1},q)$ ::=:= $(p_{0}-c)D_{0}(p_{0},p_{1},q)$
$\pi_{1}(p_{0},p_{1},q)$ ::=:= $(p_{1}-c)D_{1}(p_{0},p_{1},q).$
Following definition (2) we define naive estimators for demand and profits associated to the canonical designs:
$\widehat{GTE}_{D}(p_{0},q)$ ::=:= $\lim_{n\to\infty}\frac{D_{1}(p_{0},p_{1}^{n},q)-D_{0}(p_{0},p_{1}^{n},q)}{p_{1}^{n}-p_{0}}$
$\widehat{GTE}_{\pi}(p_{0},q)$ ::=:= $\lim_{n\to\infty}\frac{\pi_{1}(p_{0},p_{1}^{n},q)-\pi_{0}(p_{0},p_{1}^{n},q)}{p_{1}^{n}-p_{0}}$
Note that analogs of these naive estimators are frequently used in practice in estimating global treatment effects from experimental data.
@@
$\widehat{GTE}_{\pi}(p_{0},q)$ $=$ $\pi_{1y}(p_{0},p_{0},q)-\pi_{0y}(p_{0},p_{0},q)$ (9)
$=$ $D(p_{0})+(p_{0}-c)\left(D_{1y}(p_{0},p_{0},q)-D_{0y}(p_{0},p_{0},q)\right)\ ,$
@@
5 Biases: Analysis and Results
In this section we analyze biases associated to the naive estimators. Following definition (4) we define the estimation bias for demand and profits associated to the canonical designs:
$Bias_{D}(p_{0},q)$ ::=:= $GTE_{D}(p_{0})-\widehat{GTE}_{D}(p_{0},q)$
$Bias_{\pi}(p_{0},q)$ ::=:= $GTE_{\pi}(p_{0})-\widehat{GTE}_{\pi}(p_{0},q)$
@@ 5 Biases: Analysis and Results
$Bias_{D}(p_{0},q)$ $=$ $D_{0y}(p_{0},p_{0},q)+D_{1x}(p_{0},p_{0},q)$ (10)
$Bias_{\pi}(p_{0},q)$ $=$ $\pi_{0y}(p_{0},p_{0},q)+\pi_{1x}(p_{0},p_{0},q)$ (11)
$=$ $(p_{0}-c)\left(D_{0y}(p_{0},p_{0},q)+D_{1x}(p_{0},p_{0},q)\right)$
$=$ $(p_{0}-c)Bias_{D}(p_{0},q)$ (12)
As discussed in Proposition 3.1, bias increases with increasing cross-effects between treatment and control groups. More specifically, the bias is influenced by the degree to which the control (or treatment) group detracts demand from its counterpart in response to a marginal price change in the latter group.
Now, note that under Assumptions 4.2 and 4.2, both $Bias_{D}(p_{0},q)$ and $Bias_{\pi}(p_{0},q)$ are non-negative, so we obtain that the naive estimators always (weakly) underestimate $GTE$:
@@
The result follows similar cannibalization arguments developed in previous work (Johari et al., 2022). To provide intuition consider a listing-side experiment in which treatment price is larger than control price. In this case, experimental treatment demand underestimates global treatment demand because of substitution in the experiment from the treatment to the control group. Similarly, experimental control demand overestimates global control demand. Now, let’s consider a customer-side experiment where there are supply constraints, so customers compete for limited products. In such cases, the control customers may “cannibalize” supply from the treatment group, given the lower control price, leading again to an underestimation of the global treatment effect. Note that in a customer-side experiment without supply constraints, we anticipate no bias because there is no competition between the treatment and control groups.
It is important to observe that because $D^{\prime}(p)<0$ by Assumption 4, $GTE_{D}(p_{0})<0$, for all $p$. Moreover, since $\widehat{GTE}_{D}(p_{0},q)$ is an underestimation, both the estimand and the naive estimator will have the same (nonpositive) sign.
However, there is a more intriguing result for profits. Suppose that the price $p_{1}$ is greater than $p_{0}$. In this scenario, even if the demand $D(p_{1})$ at $p_{1}$ is lower than the demand $D(p_{0})$ at $p_{0}$, the increase in price may offset the decrease in demand, leading to a positive true global treatment effect in profits. However, because the naive estimator underestimates the treatment effect, it could be negative. Next, we indeed provide conditions under which the naive estimator not only underestimates the global treatment effect, but also obtains the wrong sign.
@@
Now, consider the naive estimator. Note that $\widehat{GTE}_{\pi}(p_{0},q)\leq 0$ is equivalent to $Bias_{\pi}(p_{0},q)\geq{GTE}_{\pi}(p_{0})$. That is, provided that condition $\widehat{GTE}_{\pi}(p_{0},q)\leq 0$ holds, we have a change of sign when the bias is so severe that it becomes larger than the treatment effect. Furthermore, after some manipulation of equation (9), we obtain that $\widehat{GTE}_{\pi}(p_{0},q)\leq 0$ if and only if:
@@
Note that condition (b) is a “modified markup” inequality. We have the following result:
@@
and thus the result follows because $D^{\prime}(p_{0})<0$ by Assumption 4.
The proposition states a sign change is obtained, $\widehat{GTE}_{\pi}(p_{0},q)\leq 0\leq GTE_{\pi}(p_{0})$, if and only if conditions (a) and (b) hold, that is there are appropriate lower and upper bounds on the inverse of the price elasticity. Note that we use weak inequalities in conditions (a) and (b) for a change of sign; strictly speaking of course, a change of sign can only occur if at least one of the two inequalities is strict. [^2] If exactly one of these inequalities is strict (and both hold), the canonical estimator will still instruct firms to move in a non-optimal direction; either by telling them to lower prices when they were already at the optimal price, or by suggesting they keep their price constant when they were underpricing. We suppress this edge case in the sequel in favor of simplicity of presentation. [^3] In our two-sided market model, introduced in section 7, the scenario $\widehat{GTE}_{\pi}(p_{0},q)=GTE_{\pi}(p_{0})=0$ is impossible, meaning this edge case does not exist.
To provide intuition, note that by using equation (8) we can define the experimental price elasticity of demand as:
@@
which can be understood as the experimental analogue of a markup inequality. It states that at the experimental elasticity, the current price is above the optimal monopoly price (under an appropriate concavity assumption). Hence, as expected, a change of sign occurs when the naive estimator indicates that the platform is overpricing relative to the optimal monopoly price (and therefore prefers to decrease prices), when in reality it is underpricing (and therefore should raise prices). The result has important consequences for experiment-driven price optimization: the prescription implied by the experiment (reduce price) is the opposite to the one implied by the $GTE$ (increase price). Hence, if one moves in the direction suggested by the experiment, profits decrease relative to the status-quo.
@@
We conclude with a remark regarding our estimators, our findings on changes of sign, and guidance for practitioners. Note that the estimators $\widehat{GTE}_{D}$ and $\widehat{GTE}_{\pi}$ that we defined are estimates of derivatives of the demand and profit function, respectively, as is natural given our estimands. In practice, a platform will typically directly use a difference in estimated treatment and control demand, or estimated treatment and control profit, without normalizing by the difference in prices. This can have some bearing in interpreting our results, depending on the sign of the difference between the treatment and control price.
Concretely, suppose that treatment involves raising the price relative to control. Then our results show that a change of sign can only occur if the true global treatment effect on profits is positive, but the treatment effect is estimated to be negative in the experiment. In other words, in this case, the platform risks failing to move prices in a profitable direction.
On the other hand, suppose the treatment involves lowering the price relative to control. Then our results show that a change of sign can only occur if the true global treatment effect on profits is negative, but the treatment effect is estimated to be positive in the experiment. In other words, in this case, the platform risks moving prices in a potentially unprofitable direction.

### v4

@@ 4 Revenue Management Setting
In this section we examine a traditional revenue management setting with monopoly pricing. A firm or platform offers products that, at least based on observable characteristics, are deemed homogeneous and are thus priced uniformly. The firm faces an (unknown to them) demand function given by $D(p)$ as a function of the price $p$. The firm has a constant marginal production cost, $c\geq 0$. We make the following assumption that is kept throughout the paper and will imply Assumption 3.1 for the performance metrics we study.
@@
The function $D(p)$ is differentiable and $D^{\prime}(p)<0$, for all $p\in\mathbb{R}_{+}$. At this level of abstraction, the demand function can describe a scenario with either an infinite supply or a limited supply, where a rationing mechanism determines the consumption. In the limit of a large market, even with supply constraints, we expect the demand function to be differentiable.
To provide a concrete example of a limiting regime with infinite supply, the demand function can be understood as the fraction among a large population of customers buying the product at a given price. That is, suppose $x_{i}(p)\in\{0,1\},i=1,2,\ldots$ are i.i.d. Bernoulli random variables with mean $0\leq D(p)\leq 1$; we interpret $x_{i}(p)$ as the purchasing decision of customer $i$ at price $p$. Then trivially by the law of large numbers $D(p)=\lim_{n\to\infty}\frac{1}{n}\sum_{i=1}^{n}x_{i}(p)$.
@@ 4.1 Estimands
We will consider two performance metrics: (1) demand itself, which serves as a measure of growth, and (2) the firm’s profits (or revenue, in case costs are zero). The firm’s decision variable is price, and it is seeking to gain an understanding of how changes in price impact demand (i.e., price elasticity of demand), and subsequently affect profit. To achieve this goal, the firm is interested in estimating the difference in demand or profits generated when products are priced at a new “treatment” price relative to the demand or profits generated when products are priced at the status-quo “control” price. These differences correspond to the global treatment effects of interest that we derive below. To do so, we first define the “global” profit function as:
@@
$GTE_{D}(p)$ $:=D^{\prime}(p)\ .$ (6)
$GTE_{\pi}(p)$ $:=\pi^{\prime}(p)$
$=D(p)+(p-c)D^{\prime}(p)\ .$ (7)
@@ 4.2 Experiment: Design, Model, and Estimator
Given that the demand function is unknown, we assume the firm uses price experimentation to estimate the $GTE$s of interest. As mentioned above, canonical A/B-tests can randomize on either side of the market, and our model does not preclude either possibility. That said, we do note that when conducting price experiments in practice, companies prefer to refrain from randomizing prices on the customer side to adhere to the basic principle that each customer should be offered the same price for the same product (Phillips, 2021).
@@
which represent the demand for control and treated units scaled by the control and treatment fractions, respectively. Henceforth, we refer to $D_{0}$ and $D_{1}$ as the control and treatment demand functions, respectively. Again, at this level of abstraction these demand functions could, for example, incorporate supply constraints and rationing mechanisms.
Recall that in a CR experiment treatment and control customers will observe all products priced at the treatment and control prices respectively, with the proportion of customers assigned to the treatment group being $q$. In contrast, in a LR experiment, a customer will face an array of products, a fraction $q$ of which are priced at $p_{1}$ and the rest at $p_{0}$.
We make the following assumption that implies Assumption 3.2 and that is kept throughout the paper. {assumption} The functions $D_{0}(p_{0},p_{1},q)$, and $D_{1}(p_{0},p_{1},q)$ are differentiable for all $p_{0},p_{1}\in\mathbb{R}_{+}$ and $10<q<$ 1. In addition, $p_{0},p_{1}\geq c$. Note that the assumption precludes the case of perfect substitution between control and treatment products. Under perfect substitution, a negligible undercut of the treatment relative to the control price would result in a discrete increase of the treatment demand; as a result, the demand function would not be continuous. Even though, as initially stated, the firm is selling a single type of product, we can rationalize the smoothness assumption in at least one of two ways, commonly assumed in pricing settings: (1) products can be identical in terms of observable characteristics but there may be latent horizontal differentiation; (2) there could be search frictions that do not allow customers to always buy the lowest-priced product.
We introduce the following additional assumption on the experimental demand functions which implies Assumption 3.2. {assumption} The following statements all hold:
@@
The first two statements in Assumption 4.2 are natural as they imply that control and treatment products are (imperfect) substitutes. The third assumption states that if the control and treatment prices are the same, then both the scaled control and treatment demand functions are equivalent to the market demand function; this assumption is implied by constant returns to scale. This assumption is satisfied by demand systems for which the control and treatment demand quantities are proportional to their shares in the marketplace. Indeed, this assumption will be satisfied for the model we later study in Section 7, as well for a number of other common demand systems.
@@
$\pi_{0}(p_{0},p_{1},q)$ $:=(p_{0}-c)D_{0}(p_{0},p_{1},q)\ ,$
$\pi_{1}(p_{0},p_{1},q)$ $:=(p_{1}-c)D_{1}(p_{0},p_{1},q)\ .$
Following Definition (2) we define canonical estimators for demand and profit associated with the canonical experimental designs:
$\widehat{GTE}_{D}(p_{0},q)$ $:=\lim_{n\to\infty}\frac{D_{1}(p_{0},p_{1}^{n},q)-D_{0}(p_{0},p_{1}^{n},q)}{p_{1}^{n}-p_{0}}\ ,$
$\widehat{GTE}_{\pi}(p_{0},q)$ $:=\lim_{n\to\infty}\frac{\pi_{1}(p_{0},p_{1}^{n},q)-\pi_{0}(p_{0},p_{1}^{n},q)}{p_{1}^{n}-p_{0}}\ .$
We note that analogs of these canonical estimators are frequently used in practice in estimating global treatment effects from experimental data; in particular, in our continuum model, these estimators naturally correspond to difference-in-means or Horvitz-Thompson type estimators, scaled by the price difference to estimate the target estimand (the derivative of demand or profit at the control price).
@@
$\widehat{GTE}_{\pi}(p_{0},q)$ $=\pi_{1y}(p_{0},p_{0},q)-\pi_{0y}(p_{0},p_{0},q)$
$=D(p_{0})+(p_{0}-c)\left(D_{1y}(p_{0},p_{0},q)-D_{0y}(p_{0},p_{0},q)\right)\ ,$ (9)
@@
5 Biases: Analysis and Change of Sign
In this section we analyze the biases associated with the canonical estimators introduced above. Following Definition (4), we define the estimation biases for demand and profit as:
$Bias_{D}(p_{0},q)$ $:=GTE_{D}(p_{0})-\widehat{GTE}_{D}(p_{0},q)\ ,$
$Bias_{\pi}(p_{0},q)$ $:=GTE_{\pi}(p_{0})-\widehat{GTE}_{\pi}(p_{0},q)\ .$
@@ 5 Biases: Analysis and Change of Sign
$Bias_{D}(p_{0},q)$ $=D_{0y}(p_{0},p_{0},q)+D_{1x}(p_{0},p_{0},q)\ ,$ (10)
$Bias_{\pi}(p_{0},q)$ $=\pi_{0y}(p_{0},p_{0},q)+\pi_{1x}(p_{0},p_{0},q)$
$=(p_{0}-c)\left(D_{0y}(p_{0},p_{0},q)+D_{1x}(p_{0},p_{0},q)\right)$ (11)
$=(p_{0}-c)Bias_{D}(p_{0},q)\ .$ (12)
As discussed in Proposition 3.1, bias increases as interference effects between treatment and control groups increase. More specifically, the bias is influenced by the degree to which the control (or treatment) group cannibalizes demand from its counterpart in response to a marginal price change in the latter group.
Now, note that under Assumptions 4.2 and 4.2, both $Bias_{D}(p_{0},q)$ and $Bias_{\pi}(p_{0},q)$ are non-negative, meaning the canonical estimators always (weakly) underestimate the $GTE$:
@@
This result follows similar cannibalization arguments developed in previous work (Johari et al., 2022). To provide intuition, consider an LR experiment in which the treatment price is higher than control the price. In this case, experimental treatment demand underestimates global treatment demand because of substitution in the experiment from treatment units to control units. Similarly, experimental control demand overestimates global control demand. Instead, let’s consider a CR experiment where there are supply constraints, so customers compete for limited products. In such cases, the control customers may “cannibalize” supply from the treatment group, because of their lower control price, causing an underestimation of the global treatment effect again. Note that in a CR experiment without supply constraints, we anticipate no bias because there is no competition between the treatment and control groups.
It is important to observe that because $D^{\prime}(p)<0$ by Assumption 4, $GTE_{D}(p_{0})<0$, for all $p$. Moreover, since $\widehat{GTE}_{D}(p_{0},q)$ is an underestimation, both the estimand and the canonical estimator will have the same (nonpositive) sign.
However, there is a more intriguing result for profits. Suppose that the treatment price $p_{1}$ is greater than the control price $p_{0}$. In this scenario, even if the demand $D(p_{1})$ at $p_{1}$ is lower than the demand $D(p_{0})$ at $p_{0}$, the increase in price may offset the decrease in demand, leading to a positive true global treatment effect in profits. But, because the canonical estimator underestimates the treatment effect, the profit GTE estimator could be negative. Next, we will provide conditions under which the canonical estimator not only underestimates the global treatment effect, but actually obtains the wrong sign.
@@
Now, consider the canonical estimator. Note that $\widehat{GTE}_{\pi}(p_{0},q)\leq 0$ is equivalent to $Bias_{\pi}(p_{0},q)\geq{GTE}_{\pi}(p_{0})$. That is, provided that $GTE_{\pi}(p_{0})\geq 0$ holds, we have a change of sign when the bias is so severe that it becomes larger than the treatment effect. Furthermore, after some manipulation of equation (9), we obtain that $\widehat{GTE}_{\pi}(p_{0},q)\leq 0$ if and only if:
@@
Note that condition (b) is a “modified markup” inequality. We have the following result.
@@
and thus the result follows because $D^{\prime}(p_{0})<0$ by Assumption 4. $\Box$
The proposition states a sign change is obtained, $\widehat{GTE}_{\pi}(p_{0},q)\leq 0\leq GTE_{\pi}(p_{0})$, if and only if conditions (a) and (b) hold. This implies lower and upper bounds on the inverse of the price elasticity of demand. Note that we use weak inequalities in conditions (a) and (b) for a change of sign; strictly speaking of course, a change of sign can only occur if at least one of the two inequalities is strict. [^4] If exactly one of these inequalities is strict (and both hold), the canonical estimator will still instruct firms to move in a non-optimal direction; either by telling them to lower prices when they were already at the optimal price, or by suggesting they keep their price constant when they were underpricing. We suppress this edge case in the sequel in favor of simplicity of presentation. [^5] In our two-sided market model, introduced in section 7, the scenario $\widehat{GTE}_{\pi}(p_{0},q)=GTE_{\pi}(p_{0})=0$ is impossible, meaning this edge case does not exist.
To provide intuition, note that by using equation (8) we can define the “experimental” price elasticity of demand as:
@@
which can be understood as the experimental analogue of a markup inequality. It states that the experimental elasticity suggests the current price is above the optimal monopoly price (under an appropriate concavity assumption). Hence, as expected, a change of sign occurs when the canonical estimator indicates that the platform is overpricing relative to the optimal monopoly price (and therefore prefers to decrease prices), when in reality it is underpricing (and therefore should raise prices). This result has important consequences for experiment-driven price optimization: the prescription implied an the experiment (reduce price) may be the opposite of the one implied by the $GTE$ (increase price). Therefore, if a firm moves in the direction suggested by the experiment, profits have the potential to decrease relative to the status-quo.
@@
We conclude with two remarks regarding our estimators, our findings on changes of sign, and guidance for practitioners. First, note that the estimators $\widehat{GTE}_{D}$ and $\widehat{GTE}_{\pi}$ that we defined are estimates of derivatives of the demand and profit functions, respectively, as is natural given our estimands. In practice, a platform may use a difference in estimated treatment and control demand, or estimated treatment and control profit, without normalizing by the difference in prices. This can have some bearing in interpreting our results, depending on the sign of the difference between the treatment and control price.
Concretely, suppose that treatment involves raising the price relative to control. Then our results show that a change of sign can only occur if the unnormalized global treatment effect on profits is positive, but the treatment effect is estimated to be negative in the experiment. On the other hand, suppose the treatment involves lowering the price relative to control. Then our results show that a change of sign can only occur if the unnormalized global treatment effect on profits is negative, but the treatment effect is estimated to be positive in the experiment. In either case, the platform risks either failing to move prices in a profitable direction, or worse, moving them in an unprofitable direction.
Second, we note that in our case, we have studied a model where the platform changes the price directly. In practice, another relevant platform parameter that affects profits can be a platform-wide fee. Specifically, in Appendix 12 , we present a model where a platform offers a multitude of heterogeneous goods (potentially at different prices), and earns profit by charging a percentage fee on each transaction. This is a common monetization model for online marketplaces. Viewing the fee as the variable of interest for the platform’s profit optimization, we obtain analogous results to those in this section, including the bias characterization, the possibility of a change of sign, and interpretation of the conditions for a change of sign in terms of appropriate derivatives . This extension suggests our insights in this paper also have implications for practitioners in platform settings with many heterogeneously differentiated goods, where fees are the key variable for monetization.